OpenStudy (goformit100):

The depth d(x), in feet, of a circular swimming pool at a point x feet from the edge is represented by the function below. The function d of x equals 1 plus one-eighth x] If the depth varies from 1 foot to 4 feet, which inequality describes the domain of the function? 0 ≤ x ≤ 8 0 ≤ x ≤ 24 0 ≤ x ≤ 40 1 ≤ x ≤ 8

OpenStudy (goformit100):

@stgreen

OpenStudy (goformit100):

How ?

OpenStudy (anonymous):

Well if oyu plug in 24 for x, you get 1/3 o 24=3+1=4 the outer rane of depths

OpenStudy (goformit100):

Ok...then ?

OpenStudy (anonymous):

and if oy plug in 0, you get 1/3 of 0=0+1 which is the lower range off depths that you can have

OpenStudy (goformit100):

ya got it :)